# Eka-Eval : A Comprehensive Evaluation Framework for Large Language Models in Indian Languages

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that go beyond English-centric benchmarks and address the requirements of linguistically diverse regions such as India. We present Eka-Eval, a unified and production-ready evaluation framework that integrates over 35 benchmarks, including 10 Indic-specific datasets, spanning categories like reasoning, mathematics, tool use, long-context understanding, and reading comprehension. Compared to existing Indian language evaluation tools, Eka-Eval offers broader benchmark coverage, with built-in support for distributed inference, quantization, and multi-GPU usage. Our systematic comparison positions Eka-Eval as the first end-to-end, extensible evaluation suite tailored for both global and Indic LLMs, significantly lowering the barrier to multilingual benchmarking. The framework is open-source and publicly available at https://github.com/lingo-iitgn/eka-eval and a part of ongoing Eka initiative (https://eka.soket.ai), which aims to scale up to over 100 benchmarks and establish a robust, multilingual evaluation ecosystem for LLMs.
@@
Eka-Eval : A Comprehensive Evaluation Framework for Large Language Models in Indian Languages
@@ 1 Introduction
In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023). However, these frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings. This lack of inclusivity significantly limits the effectiveness of such tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages and more than a billion native speakers.
@@
While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts, lacking integration into unified evaluation workflows. Furthermore, popular frameworks like FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption among developers and researchers operating in low-resource environments. These challenges create a need for an accessible, multilingual, and task-diverse evaluation suite.
Table 1: Comparison of LLM Evaluation Frameworks
@@
Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Multi- GPU Quantization Ready to Use
LM-Harness Medium
OpenCompass Medium
HELM Low
OpenAI Evals Low
DeepEval Medium
FreeEval Low
indic-eval High
Eka-Eval High
@@
To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs that integrates more than 35 benchmarks spanning both global (English) and Indic-language tasks. Eka-Eval is designed to be modular, easy to configure, and compatible with HuggingFace and proprietary models, offering plug-and-play usability with minimal overhead. To contextualize the capabilities of Eka-Eval, Table 1 presents a feature-level comparison of leading LLM evaluation frameworks. The comparison spans 13 dimensions, including support for long-context and tool-use benchmarks, multilingual (Indic) tasks, production readiness, and ease of use. Eka-Eval emerges as a unique blend of extensibility, multilingual support, and practical usability, combining strengths found only partially across other toolkits.
@@
Eka-Eval covers eight major evaluation categories: (i) General question answering, (ii) Mathematics and logical reasoning, (iii) Commonsense inference, (iv) World knowledge, (v) Reading comprehension, (vi) Long-context understanding, (vii) tool use and API reasoning, and (viii) Indic-specific NLP benchmarks. It includes popular benchmarks such as ARC (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that supports both global and Indic-language benchmarks, covering diverse LLM capabilities including long-context reasoning and tool use.
  • We release the framework as an open-source toolkit, designed for ease of use in real-world, low-resource, and multilingual settings to promote broad adoption and community-driven development.
Our key contributions are:
  • We propose Eka-Eval, a unified and modular evaluation framework that supports both global and Indic-language benchmarks, covering diverse LLM capabilities including long-context reasoning and tool use.
  • We release the framework as an open-source toolkit, designed for ease of use in real-world, low-resource, and multilingual settings to promote broad adoption and community-driven development.
@@ 2 Related Work
The evaluation of large language models (LLMs) has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings. This section categorizes prior work into three major strands: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations. We conclude by highlighting how Eka-Eval integrates and advances these directions.
@@
General-Purpose LLM Evaluation Frameworks Early benchmark suites such as GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages. Frameworks like HELM Liang et al. (2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models. Similarly, BIG-Bench Srivastava et al. (2022) curated over 200 tasks via crowd-sourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints. Tooling infrastructure also emerged in parallel. lm-eval-harness Gao et al. (2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals OpenAI Contributors (2023) offered YAML-based structured evaluation for internal alignment research.
@@
Recent frameworks emphasize modularity, extensibility, and accessibility. For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support. FreeEval Yu et al. (2024) introduces meta-evaluation and contamination detection. DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but has limited multilingual and low-resource customization.
@@
Specialized Capability Benchmarks To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed. Tool-use and agentic behavior are assessed in ToolBench Qin et al. (2023) and API-Bank Li et al. (2023), which evaluate model interactions with real-world APIs. Long-context reasoning is explored in InfiniteBench Zhang et al. (2024) and RULER Hsieh et al. (2024), which evaluate memory and coherence over 100K+ token sequences. Mathematical reasoning and code generation are evaluated using datasets like GSM8K Cobbe et al. (2021) and HumanEval Chen et al. (2021), respectively. While powerful, these benchmarks have limited capabilities and lack integration into unified evaluation pipelines.
@@
Multilingual and Indic Language Evaluation Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al. (2020), xP3 Muennighoff et al. (2022), and MEGA Ahuja et al. (2023), covering 40–70+ languages. However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances. FLORES Goyal et al. (2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
@@
For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages. IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability. Frameworks such as indic-eval IndicEval Contributors (2024) wrap existing tools like LightEval to support select Indic tasks. However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases.
@@
Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system. It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, while accommodating advanced model behaviors such as long-context reasoning and agentic tool use. Designed for practical deployment and large-scale experimentation, Eva-Eval includes built-in support for multi-GPU inference, quantized model evaluation, distributed execution, and customizable prompting strategies. In contrast to existing frameworks that are either limited to specific capabilities or lack multilingual extensibility, Eva-Eval offers a modular, plug-and-play architecture that enables end-to-end evaluation of LLMs across diverse languages, domains, and task configurations—bridging critical gaps in current LLM assessment infrastructures.
Figure 1: End-to-end architecture of the Eka-Eval framework.
@@ 3 Capabilities of LLM Evaluation Frameworks
We identify thirteen foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks. Some of these capabilities are described in earlier evaluation frameworks such as ….:
  • Custom Datasets: Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
  • Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
  • Custom Prompting: Provision for flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
  • Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
  • Tool Use: Evaluation of LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
  • Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
  • Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
  • Multilingual Indic: Native support for evaluation on benchmarks in Indic languages.
  • Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
  • Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
  • Multi-GPU: Built-in capability to leverage multiple GPUs for scalable and efficient inference.
  • Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit) to minimize memory requirements.
  • Ready to Use: Emphasis on ease of use with quick setup and configured tasks and models.
Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the thirteen capabilities outlined in this work. Eka-Eval demonstrates robust support across all key capabilities.
4 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability. The system is designed around three core principles: modularity for easy extension and customization, accessibility for low-barrier adoption across diverse research environments, and comprehensiveness to address the full spectrum of LLM capabilities, including underserved areas like long-context reasoning and tool use.
@@ 4.1 System Architecture
The framework follows a layered architecture consisting of four primary components: the Evaluation Engine, Benchmark Registry, Model Interface Layer, and Results Processing System, each containing specialized secondary components that handle specific functionality. This hierarchical design enables seamless integration of new benchmarks, model backends, and evaluation metrics while maintaining backwards compatibility and reproducibility.
@@ 4.1.1 Evaluation Engine
The Evaluation Engine serves as the orchestration layer, comprising three key secondary components:
Task Scheduler:
Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution as demonstrated in main_orchestrator(), which dynamically assigns evaluation tasks to available workers based on resource constraints and model requirements.
Batch Optimizer:
Implements intelligent batching strategies and supports various quantization schemes to optimize memory usage and inference speed. The optimizer automatically adjusts generation_batch_size parameters to maximize throughput while preventing out-of-memory errors, as seen in the PIQA evaluation implementation.
Distributed Coordinator:
Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library. The coordinator launches multiple worker_process instances that handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
@@ 4.1.2 Benchmark Registry
The Benchmark Registry provides a unified interface for benchmark management through two main secondary components:
Dataset Manager:
Handles diverse dataset formats and sources through the BenchmarkRegistry class, which abstracts complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
@@ 4.1.3 Model Interface Layer
The Model Interface Layer provides unified abstraction across local and API-based models through four secondary components:
Local Model Loader:
Initializes models using HuggingFace transformers with automatic device allocation, quantization support, and memory optimization. The loader handles diverse model formats while maintaining consistent interfaces.
API Client Manager:
Manages proprietary endpoints through dedicated clients (OpenAIClient, GeminiClient, ClaudeClient) that extend BaseAPIClient, providing unified request/response handling with rate limiting and authentication.
Interactive Selection Interface:
Implements get_model_selection_interface() for dynamic model discovery and selection, supporting both local model paths and API configurations with environment variable management.
Resource Manager:
Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluations and supporting smooth transitions between different model configurations.

### v2

# Eka-Eval : A Comprehensive Framework for Indian Large Language Models

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that address the requirements of linguistically diverse regions, such as India, and go beyond English-centric benchmarks. We introduce Eka-Eval, a unified evaluation framework that integrates over 35+ benchmarks (including 10 Indic benchmarks) across nine major evaluation categories. The Framework provides broader coverage than existing Indian language evaluation tools, offering 11 core capabilities through a modular architecture, seamless integration with Hugging Face and proprietary models, and plug-and-play usability. As the first end-to-end suite for scalable, multilingual LLM benchmarking, the framework combines extensive benchmarks, modular workflows, and dedicated support for low-resource Indian languages to enable inclusive assessment of LLM capabilities across diverse domains. We conducted extensive comparisons against five existing baselines, demonstrating that Eka-Eval achieves the highest participant ratings the highest participant ratings in four out of five categories.
@@
[…]
Eka-Eval : A Comprehensive Framework for Indian Large Language Models
@@ 1 Introduction
Figure 1: Architecture of Eka-Eval: A modular framework combining model inference, benchmark registry, evaluation engine, and result processing with support for caching and distributed inference.
@@
In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023). However, existing evaluation frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings (Watts et al., 2024). This lack of inclusivity significantly limits the effectiveness of these tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages encompassing over a billion native speakers.
@@
While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts and lack integration into unified evaluation workflows — requiring users to perform manual configuration to use them in real-world LLM assessment pipelines.
Furthermore, popular frameworks such as FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption by developers and researchers operating in low-resource environments. These challenges create a need for an actively maintained, community driven, multilingual, and task-diverse evaluation suite.
@@
Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Quantization
lm-eval-harness
OpenCompass
HELM
OpenAI Evals
DeepEval
FreeEval
indic-eval
Eka-Eval
Table 1: Comparison of LLM Evaluation Frameworks across eleven key capabilities, showing the comprehensive coverage of Eka-Eval.
@@
To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs, integrating over 35 benchmarks spanning both global (English) and Indic-language tasks. The overall architecture, presented in Figure 1, is designed to be modular, compatible with Hugging Face and proprietary models, it offers plug-and-play usability with minimal configuration overhead. To contextualize Eka-Eval’s capabilities, Table 1 compares leading LLM evaluation frameworks across eleven key capabilities, demonstrating its more comprehensive feature coverage relative to existing frameworks.
@@
Eka-Eval covers nine major evaluation categories: (i) Code Generation and programming; (ii) Mathematics and logical reasoning; (iii) Reading comprehension; (iv) Commonsense reasoning; (v) World knowledge; (vi) Long-context understanding; (vii) General reasoning and knowledge; (viii) Tool use and API reasoning; and (viiii) Indic-specific NLP benchmarks including ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020a), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 35 benchmarks across nine major evaluation categories, providing comprehensive assessment of both global and Indic-language LLM capabilities.
  • We provide comprehensive support for most Indian languages and conduct thorough human evaluation across five existing frameworks on four benchmark tasks, demonstrating Eka-Eval’s effectiveness and reliability.
@@ 2 Related Work
The evaluation of LLMs has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings. This section categorizes prior work into three major areas: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations. We conclude by highlighting how Eka-Eval integrates and advances these directions.
@@
General-Purpose LLM Evaluation Frameworks: Early benchmark suites such as GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages. Frameworks like HELM  (Liang et al., 2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models. Similarly, BIG-Bench (Srivastava et al., 2022) curated over 200 tasks via crowdsourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints. In parallel, tooling infrastructure also emerged. lm-eval-harness (Gao et al., 2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals (OpenAI Contributors, 2023) offered YAML-based structured evaluation for internal alignment research.
@@
Recent frameworks emphasize modularity, extensibility, and accessibility. For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support. FreeEval Yu et al. (2024) introduces meta-evaluation and contamination detection, but lacks robust multilingual and Indic language support. DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but lacks flexible multilingual and low-resource customization, and its default metrics are often too rigid.
@@
Specialized Capability Benchmarks: To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed. Tool-use and agentic behavior are assessed in ToolBench (Qin et al., 2023) and API-Bank (Li et al., 2023) respectively, which evaluate model interactions with real-world APIs.
Long-context reasoning is explored in InfiniteBench (Zhang et al., 2024) and RULER (Hsieh et al., 2024), which evaluate memory and coherence over 100K+ token sequences. Mathematical reasoning and code generation are evaluated using datasets like GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. While powerful, these benchmarks have limited scope and lack integration into unified evaluation pipelines.
@@
Multilingual and Indic Language Evaluation: Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al. (2020), xP3 Muennighoff et al. (2022), and MEGA Ahuja et al. (2023), covering 40–70 languages. However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances. FLORES Goyal et al. (2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
@@
For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages. IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability. Frameworks such as indic-eval (IndicEval Contributors, 2024) wrap existing tools, like LightEval (LightEval, 2024), to support select Indic tasks. However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases (Chen et al., 2023).
@@
Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system. It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, designed for practical deployment and large-scale experimentation.
@@ 3 Capabilities of LLM Evaluation Frameworks
We identify eleven foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks. Several of these capabilities are described in earlier evaluation frameworks such as lm-eval-harness (Gao et al., 2021), OpenCompass (OpenCompass Contributors, 2023), FreeEval (Yu et al., 2024), and DeepEval (DeepEval Contributors, 2024). These eleven capabilities are:
(1) Custom Datasets:Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
(2) Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
(3) Custom Prompting: Flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
(4) Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
(5) Tool Use: Evaluating LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
(6) Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
(7) Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
(8) Multilingual (Indic): Native support for evaluation on benchmarks in Indic languages.
(9) Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
(10) Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
(11) Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit weights) to minimize memory requirements.
To contextualize the contributions of Eka-Eval, we benchmark it against established alternatives. Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the eleven capabilities outlined in this work. Eka-Eval demonstrates robust support across all key capabilities.
4 Eka-Eval
4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability. The system is designed around three core principles: modularity for easy extension and customization; accessibility for low-barrier adoption across diverse research environments; and comprehensiveness supports a wide range of capabilities, including underserved areas such as long-context reasoning and tool use.
@@ 4.2 System Architecture
The framework follows a layered architecture consisting of four primary components:
@@ 4.2.1 Evaluation Engine
This component orchestrates all evaluation workflows:
Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution as demonstrated in main_orchestrator() - dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
Batch Optimizer: Implements intelligent batching strategies and supports various quantization schemes to optimize memory usage and inference speed. As seen in the PIQA evaluation implementation, the optimizer automatically adjusts generation_batch_size parameters to maximize throughput while preventing out-of-memory errors.
Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library. The coordinator launches multiple worker_process instances to handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
@@ 4.2.2 Benchmark Registry
Provides a unified interface for managing datasets:
Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
@@ 4.2.3 Model Interface Layer
Abstracts access to different local and API-based model backends:
Local Model Loader: Initializes transformer-based checkpoints with automatic device allocation and quantization.
API Client Manager: Manages proprietary endpoints through dedicated clients (OpenAIClient, GeminiClient, ClaudeClient) that extend BaseAPIClient, providing unified request handling with rate limiting and authentication.
Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations.
Resource Manager: Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluation runs.
Framework Setup & Config Navigation Result Export Indic Support Extensibility
[…]